OpenStudy (anonymous):

The surface area of a cube is 600 cm2. What is the area of one face of this cube?

OpenStudy (mathmale):

So the surface area is 600 cm^2? We'll need to work backward and find the length of each edge of the cube. Surface area of cube = 6x^2=600 cm^2. Find x.

OpenStudy (anonymous):

there are 6 square faces of equal area. area of each square face=600/6=100 cm^2

OpenStudy (mathmale):

Mikki: awaiting some response from you.

OpenStudy (jdoe0001):